SP-4011 Skylab: A Chronology

 

[397-400] APPENDIX 5

EXPERIMENT REPAIR AND MAINTENANCE

[From Johnson Space Center, Skylab Mission Report, Supplement 3, Flight Crew Contributions to the Skylab Mission, July 1974]

 

 

Experiment

Abnormal condition

Corrective action

Unplanned Tasks

Apollo Telescope Mount

White Light Coronagraph (S052)

Contamination on occulting disc caused bright spots on television monitor.

Contamination brushed off disc during extravehicular activity.

 

X-Ray Spectographic Telescope (S054)

Drive mechanism on aperture door failed.

Door manually opened by removing release pins during extravehicular activity.

 

Filter wheel jammed between two filter positions.

Wheel moved to open position during extravehicular activity.

 

Ultraviolet Scanning Polychromator-Spectroheliometer (S055A)

High-voltage tripoff caused by oversensitivity of detector to background radiation levels.

Main high-voltage switch configured to override for manned operation and to enable unattended operations.

 

Aperture door-ramp latch binding.

Door-ramp latch removed during extravehicular activity.

 

Extreme Ultraviolet Coronal Spectroheliograph (S082A)

Aperture door-ramp latches binding.

Door pinned open during extravehicular activity.

 

Chromospheric Extreme Ultraviolet Spectrograph (S082B)

Exposure timer operating erratically.

Replacement timer installed.

 

Cover on hydrogen-alpha 2 telescope aperture operating intermittently.

Cover pinned open during extravehicular activity.

 

Image Scope Television Monitor (S082A/B)

Monitor inoperative.

New television monitor supplied and installed successfully.

Earth Resources Experiment Package

Multispectral Scanner (S192)

Improperly seated cooler/detector.

Reseated and realigned. Alignment readings increased from 43 percent to 46 percent.

 

Incorrect prelaunch attenuator adjustments.

Special procedure uplinked and crew adjusted attenuator in flight.

 

Detector did not provide desired resolution of thermal data.

Modified thermal detector supplied and installed.

 

Microwave Radiometer/Scatterometer and Altimeter (S193)

Electrical short caused erratic antenna motion.

Antenna pinned in zero-degree pitch position during extravehicular activity, restoring operation of antenna in roll.

Biomedical

Sleep Monitoring (M133)

Cap electrodes dried out.

Rejuvenation kits used on second and third visits.

 

Mark 1 Exerciser

Mark 1 exerciser spring broke.

Spring replaced.

 

Specimen Mass Measurement Device (M074)

Electronics module failed on wardroom specimen mass measurement device.

Module replaced with electronic module from waste management compartment unit. Spare electronic module supplied and installed during next visit.

 

Experiment Repair and Maintenance-Continued

Experiment

Abnormal condition

Corrective action

Unplanned Tasks

Corollary

Nuclear Emulsion (S009)

Motor failed, preventing opening and closing of detector package.

 

Motor resupplied and replaced.

Articulated mirror system

Tilt control jammed on first visit.

Crew freed tilt adjustment gears and reported new indicator "zero" position.

Mirror surface was contaminated on second visit.

New mirror installed by third visit crew.

 

Mirror position indicator failed on third visit.

Crew positioned mirror by counting turns of control knob.

 

Ultraviolet Panorama (S183)

Film plate jammed in carrousel.

Malfunction procedure used for carrousel alignment.

Electrical failure in protective circuit, which shut off power to electronics package before exposure sequence.

 

Jumper wires connected between test connector and camera connector pins.

Contamination Measurements (T027/S073)

Photometer head in nonre tractable position due to shaft drive circuit failure.

Photometer system ejected into space, permitting anti-solar scientific airlock use in other experiments.

Student

Neutron Analysis (ED76)

Detector deactivation slide jammed.

Detector partly dismantled to free deactivation slide.

Planned Tasks

Apollo Telescope Mount

Persistence Image Scope

Fuzzy image, poorly defined bright spots and horizontal bright lines on persistence image scope.

Proper operation restored after contacts were cleaned and batteries changed.

Earth Resources Experiment Package

Multispectral Photographic Cameras (S190A)

No film motion sensing at camera station 6.

 

Camera station 6 magazine replaced with spare.

 

Dust particles on optics and film emulsion buildup on platens.

 

Cleaned with optical cleaning kits.

Tape recorder

Metal oxide buildup on recording heads.

Contamination removed from recording heads and tape rollers.

Biomedical

Sleep Monitoring (M133)

Improper operation on indicator lights.

Cable between electronic box and cap replaced.

 

Lack of telemetry data.

Two experiment cables replaced with spares.

 

Lower Body Negative Pressure (M092)

Blood pressure measuring system cuff failed.

Cuff replaced with spare.


PreviousIndexNext